Target meeting: 01Jul09,
Present: CB, TB, SG, PH, JL, KL, RN, PS, JT
Review of actions:
- JT: check whether the clean room in R2 will be available
on the 08Jul09 for the trial assembly.
- PH: Plan schedule/plan for spare fibre laying.
SG has expressed worries over personnel involved in handling delicate
fibres. Urgent that he & Willy
agree a way forward.
- SG: Liaise
with EMcC over patch panel modifications and
- PH: Obtain
spectrum and cross check with FLUKA results.
software has been obtained.
- RN: Propose
design for collar/clamp arrangement for the support of the bellows.
- PS, JL:
Write specification for generation of MICE target signal to be included in
the ISIS BPS.
- PS: Provide
signals required by ISIS as described in item 6 of the notes on the target
meeting on 17Jun09.
Under discussion with Craig. No
feedback received on go-nogo interlock.
- AN: Redraft
risk assessment and method statements for installation and removal of
target in ISIS.
- PH: will
make a target on call rota for running in September.
2. Progress on
mechanical design and fabrication:
- All drawings, including
assembly drawings, now complete.
- All parts for vacuum
testing have arrived, and are themselves being
tested. Problems found with rubber
seals – will be exchanged for copper seals. Will use this to test John Barrows’
prototype weld. Also plan to
examine section of sample weld – Eddie has provided name of contact.
- Weld of stator will not be
done for next Wednesday. Will do “dry
test” next week, and sign off design for welding. Then weld soon after.
- Polishing kit (for bearings
and shaft) has arrived at RAL.
- Shafts are with
- Issues resolved now
that expert has returned from holiday.
Machining flats is taking a long time, due to need to proceed
slowly and carefully.
- It is not possible to
reach the fine finish specified.
We will have to polish afterwards.
- 1 long shaft – welded
before grinding – (of 6 available) was damaged during grinding – it fell
apart at the weld. This could be
due to grinding too close to the weld.
It is hoped to get long shafts X-rayed at Culham
for diagnosis of extent of welded region.
- 1 short shaft (of 5)
was also damaged when flats were put on.
- JT will collect
shafts tomorrow. It is hoped
metrology can be done by the end of the week.
- A discussion of the
comparison between the two shafts followed. Due to the damage at the weld during grinding,
the short shafts may be preferable.
However we know distortion can occur during welding (though this may
be correctable). Instead of an
orbital weld, Jason suggested either stitch welds or welding through holes
in the tube (as for the stop).
However, it might be preferable to pin instead of welding. Eamonn has
made a prototype; with a single pin there is some movement, but with
crossed pins it is very solid.
Spiral pins might be better in resisting vibration. Some concern was expressed about limited
lifetime of pinning due to the hammering of the shaft, and increased wear
on the bearings if the shaft became loose or bent during use.
- A decision on welding
versus pinning should be made within a week. Tensile tests will be done at Culham to compare strengths.
have been requested for a quotation.
Andy Fox is aware of our requirements and timescale.
2.2 Stator: CB
- Parts for third stator are
available and have been checked.
Assembly has not started.
- Awaiting third housing.
to gate valve and vacuum vessel: EMcC
2.4 Assembly Test
next Wednesday: CB
- JT to chase up R2 clean
- Sheffield to bring tools,
optical block etc. Also measured magnets
3. Progress on electrical
installation etc.: SG
- DL team will be at RAL Tuesday-Thursday
- Modifications will be made
to R78 rack, patch panel etc, including control software.
- Installation at ISIS will
be done later.
- The provision of additional
signals (monitoring position of the frame) to be discussed with PS this
- The concern over fibre
installation was reiterated. To be
discussed with WS. (Fibres to be
brought from Sheffield next week.)
4. Progress on
(dynamic) mechanical FEA: SY, WL
5. Progress on
magnetic FEA: PS
electronics and DAQ: PS, JL
- Problems with
software installation caused a delay, but coding has now started.
- First writing code to
test electronics (1-2 days). Then
return to main FPGA code.
- Hardware driver
framework and EPICS channel access servers complete.
- Progressing with low
level drivers, GUI, etc.
interaction with MICE PPS and ISIS BPS: CMacW
- First pass at the plan
has been prepared (in Sheffield uploads directory). KL will post.
- Accommodation will be
booked by Sheffield team for August.
9. Next meeting
- Due to presence of
personnel at RAL for tests, this will start
at 11:00 on 8th July. Ken will
find a room.
Summary of actions:
- JT: Check whether the clean room in R2 will be available
on the 08Jul09 for the trial assembly.
- SG/KL: Arrange with WS a satisfactory method for laying spare
- SG: Liaise with EMcC over patch
panel modifications and associated wiring;
- PH: Obtain spectrum and cross check with FLUKA results.
- RN: Propose design for collar/clamp arrangement for the
support of the bellows.
- PS, JL: Write specification for generation of MICE target
signal to be included in the ISIS BPS.
- PS: Provide signals required by ISIS as described in item
6 of the notes on the target meeting on 17Jun09.
- AN: Redraft risk assessment and method statements for
installation and removal of target in ISIS.
- JT: Decision on
welding or pinning (and type of pinning) for short shafts.
- PH: Make a
target on-call rota for September.